the density of ammonia, NH3, is 7.71x10^-4 g/ml. calculate the number of miles of ammonia in a 73.1ml sample.
1
Expert's answer
2019-09-26T05:53:03-0400

The question should be to calculate no. of moles of ammonia instead of miles.

Thus for that

Density = Mass/Volume

7.71*10-4 g/ml= m/73.1

m = 0.5636 g

Moles = Sample mass/Molecular mass

Here sample mass= 0.5636g, Molecular mass of NH3 = 17g

Moles = 0.5636/17 = 0.0331